OpenStudy (anonymous):

1. Given f (x ) = x^ 2 - 4 , if f (x ) = 0, find x.

hero (hero):

f(x) = x^2 - 4 0 = x^2 - 4 0 = (x + 2)(x - 2) Finish solving for x.

OpenStudy (anonymous):

I factor? Do I factor every time I put a number in for f(x) ?

hero (hero):

You only factor if it is possible.

hero (hero):

For this setup, the appropriate alternative to factoring is completing the square if you cannot factor.